# Properties of current sheets in two-dimensional tearing-mediated magnetohydrodynamic turbulence

### v1

@@ Abstract
software: Matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), Scikit-Learn (F. Pedregosa et al., 2011), LAPS (C. Shi et al., 2024)
1 Introduction
@@
It has long been observed that the solar wind is highly turbulent, containing fluctuations on a vast span of scales (R. Bruno & V. Carbone, 2013). The solar wind turbulence is believed to be one of the major contributors to the heating and acceleration of solar wind and thus has been extensively investigated for decades (e.g. E. Marsch & C.-Y. Tu, 1990; B. Goldstein et al., 1995; R. Grappin & M. Velli, 1996; W. Matthaeus et al., 2005; C. H. Chen et al., 2012; O. Alexandrova et al., 2013; C. Shi et al., 2021; N. Sioulas et al., 2023).
@@
One of the key questions regarding solar wind turbulence is how its energy is dissipated. The widely-accepted picture is that, as the fluctuation energy cascades from large, MHD scales toward small, kinetic scales, various types of wave-particle interactions take effect, transferring energy from the turbulence to the particles (e.g. S. S. Cerri et al., 2017; J. He et al., 2019; M. M. Martinović et al., 2020; T. A. Bowen et al., 2022). In addition to these wave-particle interaction processes, intermittency naturally arises in turbulence systems, including neutral fluid (e.g. U. Frisch & G. Parisi, 1980) and plasma (e.g. L. Sorriso-Valvo et al., 2001). Intermittency refers to the sparse, irregular and strong fluctuations on top of the periodic and smooth part of the turbulence, and it breaks down the self-similarity of the turbulence. In MHD turbulence, intermittency mainly appears in the form of current sheets and vortices (M. Wan et al., 2016), and it is well-known that current sheets are usually accompanied by strong dissipation via magnetic reconnection (K. Osman et al., 2014). Therefore, intermittency may play a significant role in dissipation of MHD turbulence. In the solar wind, intermittency has been extensively observed (R. Bruno et al., 2001; W. H. Matthaeus et al., 2015; Z. Pei et al., 2016; H. Wu et al., 2023), and evidence shows that intermittency is co-located with regions of higher proton and electron temperatures in the solar wind (K. Osman et al., 2012; N. Sioulas et al., 2022b, a).
@@
The question then comes to how do we understand the generation of intermittency in MHD turbulence. In incompressible-MHD, phenomenological models were developed based on the idea of scale-dependent dynamic alignment (SDDA) (S. Boldyrev, 2005; J. Podesta et al., 2009; B. D. Chandran et al., 2015; A. Mallet et al., 2016; N. Sioulas et al., 2024a). In this scenario, the small-scale turbulence “eddies” are assumed to be anisotropic three-dimensionally. Let us write the eddy size along the background magnetic field $\bm{B_{0}}$ as $l$, the eddy size along the polarization direction of the fluctuation $\bm{\delta b}$ in the plane perpendicular to $\bm{B_{0}}$ as $\xi$, and the eddy size perpendicular to both $\bm{B_{0}}$ and $\bm{\delta b}$ as $\lambda$. In the perpendicular plane, it was assumed that the eddies are extended along the polarization direction, i.e. $\xi>\lambda$. A set of scaling relations between these lengths were derived, which were then coupled with the tearing instability theory with the assumption that the ratio $\xi/\lambda$ corresponds to the aspect ratio of the small-scale current sheets (S. Boldyrev, 2006; S. Boldyrev & N. F. Loureiro, 2017; N. F. Loureiro & S. Boldyrev, 2017; A. Mallet et al., 2017). As the scale decreases, the aspect ratio may reach a certain critical value, after which the turbulence is dominated by the recursive collapse of current sheets, i.e. the so-called “reconnection-mediated” regime, associated with a steepened power spectrum (C. Dong et al., 2022).
@@
Although both observations (e.g. J. Podesta et al., 2009; N. Sioulas et al., 2024a) and numerical simulations (e.g. J. Mason et al., 2006; J. Walker et al., 2018) have shown evidence of SDDA in plasma turbulence, there is a potential caveat in applying the SDDA theory to the reconnection in turbulence. That is, the turbulence-generated current sheets are intermittent and thus are sparse in space instead of being space-filling like the turbulent eddies. Hence, whether the anisotropic eddies directly correspond to the shape of intermittent current sheets is questionable. In this paper, we conduct and analyze a high-resolution 2D incompressible MHD simulation of balanced turbulence. The main objective is to compare the scale-dependent dynamic alignment with the properties of intermittent current sheets. We emphasize that, the classic SDDA theories are not strictly applicable to 2D turbulence due to the lack of propagation effect in 2D ($l\rightarrow+\infty$). Nonetheless, we are still able to compute the aspect ratio of the 2D eddies $\xi/\lambda$ and verify whether they are related to the aspect ratio of the current sheets.
@@
The paper is organized as follows: In Section 2, we introduce the numerical methods and describe the simulation setup. In Section 3, we present the results, including the global evolution of turbulence properties (Section 3.1), the method to identify and analyze the intermittent current sheets (Section 3.2), a case study of an evolving current sheet (Section 3.3), and the statistics of the current sheet properties (Section 3.4). In Section 4, we discuss the results with a focus on comparing SDDA with current sheet properties. In Section 5, we summarize this work.
2 Simulation setup
@@
We use the UCLA-Pseudo-Spectral (LAPS) code to conduct the simulation. LAPS is a Fourier-transform based pseudo-spectral MHD code (C. Shi et al., 2024). It has compressible and incompressible versions implemented in both 2D and 3D. Here, we use the incompressible version which allows us to compare the results with most previous theoretical models that were developed based on the assumption of incompressibility. We carry out the simulation in 2D so that high resolution can be adopted, which is essential to the development of tearing instability.
@@
are both approximately zero. Here $E_{v}=\int v^{2}$, $E_{b}=\int b^{2}$, and $E_{\pm}=\int{z^{\pm}}^{2}$ measure the total energies in velocity, magnetic field, and the two Elsässer variables $\bm{z^{\pm}}=\bm{v}\mp\bm{b}$. The root-mean-square amplitude of the initial fluctuation is $RMS(\bm{b})=0.173$. Explicit resistivity $\eta$ and viscosity $\nu$ are implemented with $\eta=\nu=5\times 10^{-7}$. Thus the (magnetic) Reynolds number is estimated to be $S_{0}=\lambda\delta b/\eta\approx 4\times 10^{4}$. We conducted tests with different amplitudes of dissipation and the values $\eta=\nu=5\times 10^{-7}$ are chosen such that the dissipation is strong enough to effectively suppress the numerical error induced by Gibbs phenomenon while not too strong so that a significant number of tearing-unstable current sheets are generated (M. Wan et al., 2013). The eddy-turnover time $\tau_{nl}=\lambda/z_{\lambda}\approx 0.7$. Thus, we run the simulation to $t=1.0$, roughly 1.4 eddy-turnover times.
Figure 3: Evolution of $J_{z}$ in the subdomain $x\in[0.2,0.7]$, $y\in[0,0.5]$. In Panel (c), the green box marks the current sheet analyzed in detail in Section 3.3 and Figure 6.
3 Results
3.1 Evolution of turbulence properties
In Figure 1, we show the time evolution of various quantities. Panel (a) shows that $\sigma_{c}$ (blue) remains roughly zero throughout the simulation, while $\sigma_{r}$ (orange) decreases rapidly at the early stage ($t\leq 0.15$), then increases a little bit and finally starts to decay slowly. The initial rapid decrease of $\sigma_{r}$ was observed in 3D simulations with similar setup (C. Shi et al., 2025). In Panel (b), we show the evolution of $E_{v}$ (blue), $E_{b}$ (orange), $E_{+}$ (green), and $E_{-}$ (red) respectively. $E_{+}$ and $E_{-}$ decay slowly due to the numerical dissipation, and so for $E_{v}$ and $E_{b}$ after $t\approx 0.3$. The early stage ($t\leq 0.15$) is featured by a rapid increase in $E_{b}$ and decrease in $E_{v}$, leading to the decrease in $\sigma_{r}$. We mark the peak of $E_{b}$ (and the valley of $E_{v}$) with the vertical dotted line. Clearly, there is a fast transfer of kinetic energy to magnetic energy at the early stage, much earlier than one eddy-turnover time. As will be shown in Figure 3, thin current sheets are formed concurrently. After $t=0.15$, reconnection becomes significant in these current sheets, transferring magnetic energy back to the kinetic energy, manifested by the slight increase in $E_{v}$. We note that, the negative residual energy is concentrated near the eddy boundaries, in regions surrounding the thin current sheets, rather than inside the current sheets where plasma jets are generated and thus featured by positive residual energy.
@@ 3.1 Evolution of turbulence properties
In Figure 2, we show the power spectra for different quantities at different time moments. We fit the spectra at $t=0.6$, when the turbulence is fully developed and the spectra are stably established over $\frac{1}{128}\leq k\leq\frac{1}{16}$, which is shown by the blue dotted-dashed lines. The fitted slopes are written in the legends. The magnetic field spectral slope is close to 5/3-5/3 and the velocity spectrum is shallower. This has been confirmed by a large number of numerical simulations (e.g. W.-C. Müller & R. Grappin, 2005; S. Boldyrev et al., 2011; C. Shi et al., 2025) and observations (e.g. J. Podesta et al., 2007; C. Chen et al., 2013; C. Shi et al., 2021). The slopes for $\bm{z^{\pm}}$ are similar and take values between 3/2-3/2 and 5/3-5/3. Figure 2 confirms that the formation of thin current sheets ($t=0.15$, Figure 1) is well before the turbulence is fully developed, since the power spectra are still evolving at $t=0.15$.
In Figure 3, we show four snapshots of $J_{z}$ in the subdomain $x\in[0.2,0.7],\,y\in[0,0.5]$. Panel (a) shows the initial condition, where $J_{z}$ is smooth (note that the color bar is in log-scale). Panel (b) shows $t=0.15$, corresponding to the vertical dotted line in Figure 1. At this moment, a number of thin current sheets are generated, but fast reconnection has not onset yet. Panel (c) shows $t=0.3$, when more current sheets are generated and the current sheets become thinner than $t=0.15$. Tearing instability has been triggered in some of the current sheets, indicated by the generation of plasmoids. One tearing-unstable current sheet is marked by the green box and this current sheet will be analyzed in detail in Section 3.3. Panel (d) shows $t=0.6$, corresponding to the vertical dashed line in Figure 1, when the turbulence is fully developed. Much more plasmoids are generated at this time compared with $t=0.3$. Figure 3 shows that the generated current sheets are located near the boundaries of eddies whose sizes are close to to the wavelengths of the initial fluctuations.
@@
at different time moments. Here $\delta\bm{b}=\bm{b}(\bm{x}+\bm{l})-\bm{b}(\bm{l})$. In general, the smaller scales correspond to larger values of Kurtosis. Interestingly, it is observed that the Kurtosis reaches maximum at $t=0.15$ and then gradually decay. Combined with Figure 1, it indicates that the increase of Kurtosis is associated with the current sheet formation process, and the later reconnection decreases the Kurtosis.
3.2 Identification of current sheets
To quantify the properties of the current sheets, it is necessary to accurately identify the current sheets. We develop an algorithm similar to that described in (V. Zhdankin et al., 2013), and we briefly describe our algorithm here. We start from finding the grid point corresponding to $\max(\left|J_{z}\right|)$ throughout the whole domain, excluding regions already marked as current sheets. Then we run a recursive, depth-first search algorithm to check the four neighboring grid points. The algorithm returns when all the neighboring points have $\left|J_{z}\right|\leq 0.1\max(\left|J_{z}\right|)$, and we mark the continuous set of data points as one single current sheet. This current-sheet search is repeated until all the current sheets with peak-$\left|J_{z}\right|$ larger than $0.1\max_{global}(\left|J_{z}\right|)$ are identified, where $\max_{global}(\left|J_{z}\right|)$ is the maximum $\left|J_{z}\right|$ in the whole simulation domain. In Figure 5(a), we show the same plot as Figure 3(c) and mark the identified current sheets in yellow ($J_{z}>0$) and green ($J_{z}<0$). It shows that the algorithm works well, and most of the identified current sheets exist between the interacting large-scale “eddies”, i.e. near the eddy boundaries. However, we note that, as the turbulence develops, there are many regions with strong current density but are “broken” instead of being coherent structures. Thus, we further refine the current sheet lists by discarding regions with less than 20 data points.
@@ 3.2 Identification of current sheets
Figure 6: Time evolution of a single current sheet within the domain marked by the green box in Figure 3. In each panel, top plot shows $J_{z}$ and bottom plot shows $\omega_{z}$. In panels (a) & (b), the green shade is the automatically identified current sheet region using the algorithm described in Section 3.2. The white and red lines correspond to the segments of the current sheet given by the recursive PCA algorithm described in Section 3.3.
3.3 Case study of a current sheet
We aim to accurately quantify the geometry of these current sheets, including their lengths and thicknesses. Therefore, a principal component analysis (PCA) based algorithm is developed, with the aid of PCA functions implemented in the Python package Scikit-Learn (F. Pedregosa et al., 2011).

### v2

@@ Abstract
software: Matplotlib (Hunter2007Matplotlib), NumPy (harris2020array), Scikit-Learn (scikit-learn), LAPS (shi2024laps)
I Introduction
@@
It has long been observed that the solar wind is highly turbulent, containing fluctuations on a vast span of scales (bruno2013solar). The solar wind turbulence is believed to be one of the major contributors to the heating and acceleration of solar wind and thus has been extensively investigated for decades (e.g. marsch1990spectral; goldstein1995properties; grappin1996waves; matthaeus2005spatial; chen2012three; alexandrova2013solar; shi2021alfvenic; sioulas2023magnetic).
@@
One of the key questions regarding solar wind turbulence is how its energy is dissipated. The widely-accepted picture is that, as the fluctuation energy cascades from large, MHD scales toward small, kinetic scales, various types of wave-particle interactions take effect, transferring energy from the turbulence to the particles (e.g. cerri2017plasma; he2019direct; martinovic2020enhancement; bowen2022situ). In addition to these wave-particle interaction processes, intermittency naturally arises in turbulence systems, including neutral fluid (e.g. frisch1980fully) and plasma (e.g. sorriso2001intermittency). Intermittency refers to the sparse, irregular and strong fluctuations on top of the periodic and smooth part of the turbulence, and it breaks down the self-similarity of the turbulence. In MHD turbulence, intermittency mainly appears in the form of current sheets and vortices (wan2016intermittency), and it is well-known that current sheets are usually accompanied by strong dissipation via magnetic reconnection (osman2014magnetic). Therefore, intermittency may play a significant role in dissipation of MHD turbulence. In the solar wind, intermittency has been extensively observed (bruno2001identifying; matthaeus2015intermittency; pei2016influence; wu2023intermittency), and evidence shows that intermittency is co-located with regions of higher proton and electron temperatures in the solar wind (osman2012intermittency; sioulas2022statistical; sioulas2022preferential).
@@
The question then comes to how do we understand the generation of intermittency in MHD turbulence. In incompressible-MHD, phenomenological models were developed based on the idea of scale-dependent dynamic alignment (SDDA) (boldyrev2005spectrum; podesta2009scale; chandran2015intermittency; mallet2016measures). In this scenario, the small-scale turbulence “eddies” are assumed to be anisotropic three-dimensionally. Let us write the eddy size along the background magnetic field $\bm{B_{0}}$ as $l$, the eddy size along the polarization direction of the fluctuation $\bm{\delta b}$ in the plane perpendicular to $\bm{B_{0}}$ as $\xi$, and the eddy size perpendicular to both $\bm{B_{0}}$ and $\bm{\delta b}$ as $\lambda$. In the perpendicular plane, it was assumed that the eddies are extended along the polarization direction, i.e. $\xi>\lambda$. A set of scaling relations between these lengths were derived, which were then coupled with the tearing instability theory with the assumption that the ratio $\xi/\lambda$ corresponds to the aspect ratio of the small-scale current sheets (boldyrev2006spectrum; boldyrev2017magnetohydrodynamic; loureiro2017role; mallet2017disruption). As the scale decreases, the aspect ratio may reach a certain critical value, after which the turbulence is dominated by the recursive collapse of current sheets, i.e. the so-called “reconnection-mediated” regime, associated with a steepened power spectrum (dong2022reconnection).
@@
Although both observations (e.g. podesta2009scale; sioulas2024scale) and numerical simulations (e.g. mason2006dynamic; walker2018influence) have shown evidence of SDDA in plasma turbulence, there is a potential caveat in applying the SDDA theory to the reconnection in turbulence. That is, the turbulence-generated current sheets are intermittent and thus are sparse in space instead of being space-filling like the turbulent eddies. Hence, whether the anisotropic eddies directly correspond to the shape of intermittent current sheets is questionable. In this paper, we conduct and analyze a high-resolution 2D incompressible MHD simulation of balanced turbulence. The main objective is to compare the scale-dependent dynamic alignment with the properties of intermittent current sheets. We emphasize that, the classic SDDA theories are not strictly applicable to 2D turbulence due to the lack of propagation effect in 2D ($l\rightarrow+\infty$). Nonetheless, we are still able to compute the aspect ratio of the 2D eddies $\xi/\lambda$ and verify whether they are related to the aspect ratio of the current sheets.
@@
The paper is organized as follows: In Section II, we introduce the numerical methods and describe the simulation setup. In Section III, we present the results, including the global evolution of turbulence properties (Section III.1), the method to identify and analyze the intermittent current sheets (Section III.2), a case study of an evolving current sheet (Section III.3), and the statistics of the current sheet properties (Section III.4). In Section IV, we discuss the results with a focus on comparing SDDA with current sheet properties. In Section V, we summarize this work.
II Simulation setup
@@
We use the UCLA-Pseudo-Spectral (LAPS) code to conduct the simulation. LAPS is a Fourier-transform based pseudo-spectral MHD code (shi2024laps). It has compressible and incompressible versions implemented in both 2D and 3D. Here, we use the incompressible version which allows us to compare the results with most previous theoretical models that were developed based on the assumption of incompressibility. We carry out the simulation in 2D so that high resolution can be adopted, which is essential to the development of tearing instability.
@@
are both approximately zero. Here $E_{v}=\int v^{2}$, $E_{b}=\int b^{2}$, and $E_{\pm}=\int{z^{\pm}}^{2}$ measure the total energies in velocity, magnetic field, and the two Elsässer variables $\bm{z^{\pm}}=\bm{v}\mp\bm{b}$. The root-mean-square amplitude of the initial fluctuation is $RMS(\bm{b})=0.173$. Explicit resistivity $\eta$ and viscosity $\nu$ are implemented with $\eta=\nu=5\times 10^{-7}$. Thus the (magnetic) Reynolds number is estimated to be $S_{0}=\lambda\delta b/\eta\approx 4\times 10^{4}$. We conducted tests with different amplitudes of dissipation and the values $\eta=\nu=5\times 10^{-7}$ are chosen such that the dissipation is strong enough to effectively suppress the numerical error induced by Gibbs phenomenon while not too strong so that a significant number of tearing-unstable current sheets are generated (wan2013generation). The eddy-turnover time $\tau_{nl}=\lambda/z_{\lambda}\approx 0.7$. Thus, we run the simulation to $t=1.0$, roughly 1.4 eddy-turnover times.
Figure 3: Evolution of $J_{z}$ in the subdomain $x\in[0.2,0.7]$, $y\in[0,0.5]$. In Panel (c), the green box marks the current sheet analyzed in detail in Section III.3 and Figure 6.
III Results
III.1 Evolution of turbulence properties
In Figure 1, we show the time evolution of various quantities. Panel (a) shows that $\sigma_{c}$ (blue) remains roughly zero throughout the simulation, while $\sigma_{r}$ (orange) decreases rapidly at the early stage ($t\leq 0.15$), then increases a little bit and finally starts to decay slowly. The initial rapid decrease of $\sigma_{r}$ was observed in 3D simulations with similar setup (shi2025evolution). In Panel (b), we show the evolution of $E_{v}$ (blue), $E_{b}$ (orange), $E_{+}$ (green), and $E_{-}$ (red) respectively. $E_{+}$ and $E_{-}$ decay slowly due to the numerical dissipation, and so for $E_{v}$ and $E_{b}$ after $t\approx 0.3$. The early stage ($t\leq 0.15$) is featured by a rapid increase in $E_{b}$ and decrease in $E_{v}$, leading to the decrease in $\sigma_{r}$. We mark the peak of $E_{b}$ (and the valley of $E_{v}$) with the vertical dotted line. Clearly, there is a fast transfer of kinetic energy to magnetic energy at the early stage, much earlier than one eddy-turnover time. As will be shown in Figure 3, thin current sheets are formed concurrently. After $t=0.15$, reconnection becomes significant in these current sheets, transferring magnetic energy back to the kinetic energy, manifested by the slight increase in $E_{v}$. We note that, the negative residual energy is concentrated near the eddy boundaries, in regions surrounding the thin current sheets, rather than inside the current sheets where plasma jets are generated and thus featured by positive residual energy.
@@ III.1 Evolution of turbulence properties
In Figure 2, we show the power spectra for different quantities at different time moments. We fit the spectra at $t=0.6$, when the turbulence is fully developed and the spectra are stably established over $\frac{1}{128}\leq k\leq\frac{1}{16}$, which is shown by the blue dotted-dashed lines. The fitted slopes are written in the legends. The magnetic field spectral slope is close to 5/3-5/3 and the velocity spectrum is shallower. This has been confirmed by a large number of numerical simulations (e.g. muller2005spectral; boldyrev2011spectral; papini2019can; shi2025evolution) and observations (e.g. podesta2007spectral; chen2013residual; shi2021alfvenic). The slopes for $\bm{z^{\pm}}$ are similar and take values between 3/2-3/2 and 5/3-5/3. Figure 2 confirms that the formation of thin current sheets ($t=0.15$, Figure 1) is well before the turbulence is fully developed, since the power spectra are still evolving at $t=0.15$.
In Figure 3, we show four snapshots of $J_{z}$ in the subdomain $x\in[0.2,0.7],\,y\in[0,0.5]$. Panel (a) shows the initial condition, where $J_{z}$ is smooth (note that the color bar is in log-scale). Panel (b) shows $t=0.15$, corresponding to the vertical dotted line in Figure 1. At this moment, a number of thin current sheets are generated, but fast reconnection has not onset yet. Panel (c) shows $t=0.3$, when more current sheets are generated and the current sheets become thinner than $t=0.15$. Tearing instability has been triggered in some of the current sheets, indicated by the generation of plasmoids. One tearing-unstable current sheet is marked by the green box and this current sheet will be analyzed in detail in Section III.3. Panel (d) shows $t=0.6$, corresponding to the vertical dashed line in Figure 1, when the turbulence is fully developed. Much more plasmoids are generated at this time compared with $t=0.3$. Figure 3 shows that the generated current sheets are located near the boundaries of eddies whose sizes are close to to the wavelengths of the initial fluctuations.
@@
at different time moments. Here $\delta\bm{b}=\bm{b}(\bm{x}+\bm{l})-\bm{b}(\bm{l})$. In general, the smaller scales correspond to larger values of Kurtosis. Interestingly, it is observed that the Kurtosis reaches maximum at $t=0.15$ and then gradually decay. Combined with Figure 1, it indicates that the increase of Kurtosis is associated with the current sheet formation process, and the later reconnection results in a decrease of the Kurtosis.
III.2 Identification of current sheets
To quantify the properties of the current sheets, it is necessary to accurately identify the current sheets. We develop an algorithm similar to that described in (zhdankin2013statistical), and we briefly describe our algorithm here. We start from finding the grid point corresponding to $\max(\left|J_{z}\right|)$ throughout the whole domain, excluding regions already marked as current sheets. Then we run a recursive, depth-first search algorithm to check the four neighboring grid points. The algorithm returns when all the neighboring points have $\left|J_{z}\right|\leq 0.1\max(\left|J_{z}\right|)$, and we mark the continuous set of data points as one single current sheet. This current-sheet search is repeated until all the current sheets with peak-$\left|J_{z}\right|$ larger than $0.1\max_{global}(\left|J_{z}\right|)$ are identified, where $\max_{global}(\left|J_{z}\right|)$ is the maximum $\left|J_{z}\right|$ in the whole simulation domain. In Figure 5(a), we show the same plot as Figure 3(c) and mark the identified current sheets in yellow ($J_{z}>0$) and green ($J_{z}<0$). It shows that the algorithm works well, and most of the identified current sheets exist between the interacting large-scale “eddies”, i.e. near the eddy boundaries, as pointed out by papini2019can in their Hall-MHD and hybrid Particle-in-Cell simulations. However, we note that, as the turbulence develops, there are many regions with strong current density but are “broken” instead of being coherent structures. Thus, we further refine the current sheet lists by discarding regions with less than 20 data points.
@@ III.2 Identification of current sheets
Figure 6: Time evolution of a single current sheet within the domain marked by the green box in Figure 3. In each panel, top plot shows $J_{z}$ and bottom plot shows $\omega_{z}$. In panels (a) & (b), the green shade is the automatically identified current sheet region using the algorithm described in Section III.2. The white and red lines correspond to the segments of the current sheet given by the recursive PCA algorithm described in Section III.3.
III.3 Case study of a current sheet
We aim to accurately quantify the geometry of these current sheets, including their lengths and thicknesses. Therefore, a principal component analysis (PCA) based algorithm is developed, with the aid of PCA functions implemented in the Python package Scikit-Learn (scikit-learn).